# Photonic-Electronic Integrated Circuits for High-Performance Computing and AI Accelerator

### v1

@@ Abstract
In recent decades, the demand for computational power has surged, particularly with the rapid expansion of artificial intelligence (AI). As we navigate the post-Moore’s law era, the limitations of traditional electrical digital computing, including process bottlenecks and power consumption issue, are propelling the search for alternative computing paradigms. Among various emerging technologies, integrated photonics stands out as a promising solution for the next generation of high-performance computing due to the inherent advantages of light, such as low latency, high bandwidth, and unique multiplexing techniques. Furthermore, the progress in photonic integrated circuits (PICs), which are equipped with abundant photoelectronic components, positions photonic-electronic integrated circuits as a viable solution for high-performance computing and as hardware AI accelerators. In this review, we survey recent advancements in both PIC-based digital and analog computing for AI, exploring the principal benefits and obstacles of implementation. Additionally, we propose a comprehensive analysis of photonic AI from the perspectives of hardware implementation, accelerator architecture, and software-hardware co-design. In the end, acknowledging the existing challenges, we underscore potential strategies for overcoming these issues and offer insights into the future drivers for optical computing.
@@ I Introduction
Since its inception in 1965, Moore’s Law has played a pivotal role in the evolution of semiconductor technology, driving the ongoing reduction of transistor feature sizes and enhanced performance of electronic devices[1]. In recent years, the proliferation of deep neural networks (DNNs) has been remarkable, exerting increasing influence across various aspects of our lives, including image recognition [2, 3, 4, 5], natural language processing [6, 7], autonomous driving [8], and medical diagnosis [9, 10]. One recent notable example is the emergence of large language models (LLMs) such as GPT (Generative Pre-trained Transformer). These models exhibit human-level intelligence and have revolutionized a wide range of applications, from sophisticated chatbots to advanced text analysis tools. However, the advancements of DNNs are driven by rapidly escalating model sizes and data volumes, necessitating massively growing computational demands. For instance, the GPT-3 model developed by OpenAI, which contains around 175 billion parameters, requires approximately 288 years for training with a single NVIDIA V100 GPU [11, 12]. Hence, there is a pressing need to develop processing units capable of performing high-speed, energy-efficient computing to meet the computational demands of AI and the big data era. Nevertheless, as process nodes progress to several-nanometer scales, the industry faces the inherent physical limits of materials[13, 14]. A primary concern is the surge in power consumption with clock frequencies reaching up to a few gigahertz, leading to overwhelming heat generation[14, 15]. Furthermore, at these diminutive scales, quantum uncertainties start to dominate electron behavior, resulting in increased transistor errors and reduced reliability [16]. Since then, both industry and academia are actively exploring alternative avenues from novel materials[17, 18], architectures[2, 19], and the investigation of new computational paradigms to meet the surging demand of high-performance computing.
@@
Among the emerging technologies, integrated photonics are promising candidates for next-generation computation by addressing the aforementioned bottlenecks of their electrical counterparts. First, optical signals travel at the speed of light within optical waveguides, surpassing electron-based transit through transistors with multiple fanouts by 1-2 orders of magnitude[20]. The delay and loss in waveguides are primarily determined by the optical path length. Additionally, a series of high-speed and energy-efficient (operating on the order of sub-picojoule per bit) devices for optical computing have been developed[21, 22, 23]. Notably, the power consumption of transistor-based electrical circuits exhibits a cubic relationship with the clock frequency $f$[24], whereas photonics-electronic platform scale only linearly with $f$[25], effectively unleashes the frequency constraints associated with the power wall issue. Furthermore, as bosons, photons do not conform to the Pauli exclusion principle, allowing for the utilization of unique multiplexing techniques such as wavelength division multiplexing (WDM) and polarization division multiplexing (PDM), which further increase the overall aggregate bandwidth. Moreover, compared to another photonic computing scheme in the form of free-space diffraction [26, 27], integrated photonics offers much greater compactness for high integration. The advancement of silicon photonics has enabled the implementation of optical computing on low-cost photonic integrated circuits (PICs) with high integration density, leveraging CMOS-compatible silicon manufacturing techniques As an increasing number of foundries develop their validated process design kits (PDKs), the integrated photonics industry is progressively moving towards standardization similar to that of the fabless semiconductor industry [28]. This trend not only improves accessibility for designers and users but also offers more reliable performance guarantees.
@@
Integrated photonics has shown itself as a promising platform for AI accelerators, benefiting from its inherent attributes of high parallelism, low latency, and power consumption. In the last decade, a diverse range of PIC-based optical neural networks (ONNs) that implement multilayer perceptions (MLPs)[29, 30], convolutional neural networks (CNNs)[27, 31, 32, 33], spiking neural networks (SNNs)[34, 35], etc., have been reported, demonstrating remarkable performance on machine learning tasks. The fundamental operations of neural networks, involving data transfers and tensor operations, are achieved through the combination of passive optical devices and high-performance active photonic-electronic components[36, 37]. Specifically, optical signals can be modulated by electrical signals and “multiplied” in accordance with the transmission function of PIC. The hybrid photonic-electronic platform combines the adaptability of electronic control with the high-speed capabilities of optical computing. Recently, cutting-edge optical processing units have been reported with a matrix processing speed of 3.8 trillion operations per second (TOPS) via time-wavelength multiplexing[32], while other works demonstrated ultra-low power consumption on the order of sub-femtojoules per bit[23, 38].
@@
While integrated photonics offers new opportunities for high-performance computing, existing photonic-electronic computing systems still encounter several practical challenges, such as:
@@
  • Inherent challenges in PICs, such as the on-chip implementation of nonlinearity for ONNs and system robustness against noise and crosstalk, require careful consideration and refinement in both hardware design and software coordination.
This review focuses on recent progress in photonic-electronic integrated circuits for computing. Spanning from digital computing to analog AI accelerators, this paper is structured as follows. Section II begins with an overview of the fundamental blocks in PIC-based digital computing, followed by a survey of recent highlight research ranging from the implementation of logic gates to fully functional photonic processing units. Section III focuses on the implementation of ONNs, covering aspects of photonic tensor cores, nonlinearity, and hardware-aware training strategies. Beyond a review from the device and circuit level, Sections IV and V provide a comprehensive analysis of recent photonic AI efforts from the perspectives of AI accelerator architectures and software-hardware co-design, respectively. The review culminates with Section VI, which offers an outlook on PIC-based optical computing and provides a summarizing conclusion.
II Survey of Optical Digital Computing on PICs
@@ II-A Optical Logic Gates
In the digital domain, both input and output are binary, and the resolution is defined by the number of bits and remains unaffected by the circuit size. A range of building blocks for optical digital computing on integrated photonic platforms, such as optical switches [38], modulators [23, 39], interconnects and photodetectors [40, 41, 42, 43], have been experimentally demonstrated. Basic logic operations (NOT, AND, OR, XOR, etc.), which are fundamental elements of digital systems, have been developed with diverse mechanisms. Among these, electro-optic logic, also known as optical-directed logic, has been widely investigated by many research groups as well as foundries. As shown in Fig. 1.a, all E-O devices in the functional block, such as Mach-Zehnder Interferometer (MZI), microring resonator (MRR), and microdisk, are simultaneously configured by electrical signals. When light traverses the block, optical signals are modulated to execute logic operations in accordance with PIC design and then propagated downstream or detected by monitors to read out the results. An important feature of the E-O digital logic is that each device is controlled by independent electrical input simultaneously and that signals are transmitted via light without the limitation of RC time constant and delay accumulations inherent in electrical systems. In other words, E-O logic merges the convenience and flexibility of electrical control with the high-speed capabilities of optical computing. Fig. 1.b-c show examples of E-O logic gates using two cascaded MRRs to perform 2-input AND/NAND and OR/NOR operations [44]. The proposed optical logic gate leverages the transmission characteristic of add-drop MRRs, which work as optical switches to implement logic operation and generate complementary outputs from the add and drop port. Additionally, by configuring the resonant states of MRRs with ”0” operands through the selection of continuous wave (CW) inputs at on-resonance or off-resonance wavelengths, the identical photonic circuit can perform various logic functions. With consistent logic but varying PIC topologies, XOR and XNOR gates featuring a crossbar structure also have been demonstrated  [45] (Fig. 1.d). Operating with a similar mechanism, MZI-based optical gates are also extensively developed [46, 47].
@@
Figure 1: Implementations of electro-optic logic gates. (a) Schematic diagram of E-O logic composed of passive and active optical components. During each clock cycle, electrical inputs are used to configure the logic circuit, while light carries out logic operations based on the transmission characteristic of the logic block. (b)-(d) Schematic of MRR-based AND/NAND, OR/NOR, and XOR/XNOR gates, as proposed in Ref [44, 45]. In the diagram, the dotted line and the solid line represent MRRs functioning in the “block/pass” and “pass/block” modes, respectively. These configurations correspond to the outputs of ”0” and ”1” at the through port, given a logic ”0” as the electrical input.
@@
In addition to E-O modulation, all-optical logic gate devices also attracted attention. These devices have been experimentally demonstrated using various techniques, including photonic crystals [48, 49], surface plasmon polariton (SPP) [50], nanowire networks [51], and slot waveguides [52, 53]. A comprehensive classification, explanation of mechanisms, and comparative analysis of all-optical logic gates have been detailed in the prior review [54, 55, 56]. Compared to E-O logic, all-optical logic offers the potential higher operation speed and bandwidth and eliminates the extra energy consumption associated with O-E-O conversion. However, the implementation of all-optical logic confronts several practical challenges. Firstly, the complexity and requirement of design and fabrication (e.g., the transmission characteristic of photonic crystal is highly sensitive to lattice constants), lead to inherent instability and a low contrast ratio between logical states in all-optical gates [53]. Secondly, while all-optical gates can be more energy-efficient in signal processing, they often require higher optical signal power to compensate for higher losses or to induce the necessary non-linear effects for switching [57]. Furthermore, the limited functionality and scalability of all-optical logic, along with higher cost, also restrict its scalability and widespread applicability compared to E-O logic.
II-B Combinational Logic and Reconfigurable PIC
@@
In digital circuits, the output of a combinational logic unit is determined only by the current input combination, without dependence on previous or future states. Similarly, the implementation of optical combinational logic could begin with extracting the logical expression from its truth table, followed by designing the corresponding PIC based on the simplified expression. Implementation can rely on assembling fundamental logic gates or leveraging the unique characteristics of optical components or multiplexing techniques for fewer devices and compact layouts. Employing these strategies, diverse optical combinational logic units, including but not limited to adders [58], comparators [59], encoders [60, 61] and decoders [62], have been reported.
@@
The aforementioned optical logic gates and units, tailored for specific tasks, are constrained by a fixed or limited logic representation. The inherent limitation not only complicates the development process but also increases the cost. To address this challenge, reconfigurable PICs offer a promising solution by programming the operational states of optical switches within a pre-designed framework using additional signals. The input operands and the reconfiguration signal can be independently managed by two separate control units within the modulator, such as the dual arms of MZIs. Generally, reconfiguration signals do not require high-speed modulation to the same extent as the input signals used as logic operands. Qiu et al. proposed a reconfigurable logic unit based on the MRRs embedded with two modulation mechanisms  [63]. As is shown in Fig. 2.a, the logic operand is modulated at high speeds via the p-i-n junction operating in the carrier injection mode, while the resonant state of MRR can be reconfigured using the microheater. Furthermore, multi-operand modulators and non-volatile devices are also promising candidates for reconfigurable PICs [64, 65].
@@
$Y=y_{1}+y_{2}+\ldots+y_{m},\text{ where }y_{i}=\prod_{k=1}^{n}X_{k}\text{ or }\overline{X_{k}}$ (1)
@@
In this expression, $\overline{X_{k}}$ denotes the complement logic of input $X_{k}$. Using reconfigurable optical switches, the architecture illustrated in Fig. 2.b-c theoretically can implement arbitrary logic functions conforming to the expression format presented in Eq. (1). The product term $y_{i}$, i.e., logical AND operation, can be implemented by $n$ serially connected reconfigurable optical switches along single bus waveguides. This block yields logic ”1” output only when all switches are in the ”pass” state. Reconfiguration signals $R_{i}$ are used to determine whether the corresponding operand contributes complement logic to $y_{i}$. In contrast to logic computing within the electrical domain, the OR operation for optical signals can be directly implemented using a combiner and detected by the photodetector. It is important to recognize that when product terms are represented by the same wavelength, the amalgamation may result in logical errors due to coherent interference. This issue can be avoided by either using distinct wavelengths for each branch [63], or by equipping each branch with photodetectors individually [66]. Besides, given that photodetectors operate as current sources, a straightforward parallel configuration could achieve electrical OR logic without introducing additional delays. While this strategy offers a universal solution for optical combinational logic, the limited scalability restricts its practical applicability in scenarios involving numerous operands. For an $n$-operands system, the complexity of this processing unit escalates as $O(n\cdot 2^{n})$, indicating an exponential increase in the requisite number of switches with $n$. Furthermore, issues related to power consumption and accumulated losses also need to be considered.
@@
Figure 2: Reconfigurable PICs for arbitrary combinational logic. (a). Optical micrograph and cross-sectional diagram of a reconfigurable MRR featuring two modulation mechanisms. The p-i-n junction is the applied by RF signal for input encoding, while the microheater is connected to the low-speed DC signal for resonance mode reconfiguration [63]. (b)-(c) Schematic of the PIC architecture enables the implementation of arbitrary combinational logic expressions based on reconfigurable optical switch (ROS).
@@
In the realm of electrical digital design, electronic design automation (EDA) tools assist designers by automatically generating and optimizing circuits from logic expressions, as well as auto-placing and routing in physical design. With the advancement of PIC-based computing, logic synthesis methods specified for PIC have been proposed, enabling large-scale design automation and optimization of compact PICs for digital computing [67, 68, 69].
II-C Toward Fully-functional EPALU

### v2

@@ Abstract
In recent decades, the demand for computational power has surged, particularly with the rapid expansion of artificial intelligence (AI). As we navigate the post-Moore’s law era, the limitations of traditional electrical digital computing, including process bottlenecks and power consumption issues, are propelling the search for alternative computing paradigms. Among various emerging technologies, integrated photonics stands out as a promising solution for next-generation high-performance computing, thanks to the inherent advantages of light, such as low latency, high bandwidth, and unique multiplexing techniques. Furthermore, the progress in photonic integrated circuits (PICs), which are equipped with abundant photoelectronic components, positions photonic-electronic integrated circuits as a viable solution for high-performance computing and hardware AI accelerators. In this review, we survey recent advancements in both PIC-based digital and analog computing for AI, exploring the principal benefits and obstacles of implementation. Additionally, we propose a comprehensive analysis of photonic AI from the perspectives of hardware implementation, accelerator architecture, and software-hardware co-design. In the end, acknowledging the existing challenges, we underscore potential strategies for overcoming these issues and offer insights into the future drivers for optical computing.
@@ I Introduction
As the semiconductor industry advances to process nodes below 3 nanometers, it increasingly encounters inherent physical limitations of both devices and materials [1, 2]. A primary concern is the surge in power consumption as clock frequencies reach gigahertz levels, leading to overwhelming heat generation [2, 3]. Furthermore, at these diminutive scales, quantum uncertainties begin to dominate electron behavior, resulting in increased transistor errors and reduced reliability. Additionally, AI has made remarkable strides in recent years, exerting a growing influence on various aspects of our lives, such as image recognition [4, 5, 6, 7], natural language processing [8, 9], autonomous driving [10], and medical diagnosis [11, 12], which have further increased societal demand for computational power. One notable example is the emergence of large language models (LLMs) such as GPT (Generative Pre-trained Transformer). These models exhibit human-level intelligence and have revolutionized a wide range of applications, from sophisticated chatbots to advanced text analysis tools. However, the advancements of deep neural networks (DNNs) are driven by rapidly increasing model sizes and data volumes, which necessitate significantly expanding computational demands. For instance, the GPT-3 model developed by OpenAI, which contains around 175 billion parameters, requires 14.8 days for training using a cluster of around 10,000 NVIDIA V100 GPUs, with an estimated energy consumption of 1287 MWh [13, 14, 15]. Hence, in the post-Moore’s Law era, traditional electronic computing architectures, designed to execute sequential, digital programs, are inadequate to meet the surging demand for high-performance computing and AI tasks. There is a pressing need to develop processing units capable of performing high-speed, energy-efficient computing. In response, both industry and academia are actively exploring alternative avenues from novel materials [16, 17], architectures [4, 18], to the investigation of new computational paradigms.
@@
Among the emerging technologies, integrated photonics is a promising candidate for next-generation computation that can overcome the bottlenecks of their electrical counterparts. First, the speed of optical signals travel within optical waveguides surpasses that of electron-based transit through transistors with multiple fanouts by 1-2 orders of magnitude [19]. The delay and loss in waveguides are primarily determined by the optical path length. Additionally, a series of high-speed and energy-efficient (operating on the order of sub-picojoule per bit) devices for optical computing have been developed[20, 21, 22]. Notably, the power consumption of transistor-based electrical circuits exhibits a cubic relationship with the clock frequency $f$[23], whereas photonics-electronic platforms scale only linearly with $f$[24], effectively relaxing the frequency constraints associated with the power wall issue. Furthermore, as bosons, photons do not conform to the Pauli exclusion principle, allowing for the utilization of unique multiplexing techniques such as wavelength division multiplexing (WDM), which further increases the overall bandwidth. Moreover, compared to another photonic computing scheme in the form of free-space diffraction [25, 26], integrated photonics offers superior compactness for higher-level integration. The advancement of silicon photonics has enabled the implementation of optical computing on low-cost PICs with high integration density, leveraging CMOS-compatible silicon manufacturing techniques. As an increasing number of foundries develop their validated process design kits (PDKs), the integrated photonics industry is progressively moving towards standardization similar to that of the fabless semiconductor industry [27]. This trend not only improves accessibility for designers and users but also offers more reliable performance.
@@
Integrated photonics has emerged as a promising platform for AI accelerators, benefiting from its inherent attributes of high parallelism, low latency, and low power consumption. In the last decade, a diverse range of PIC-based optical neural networks (ONNs) that implement multilayer perceptrons (MLPs)[28], convolutional neural networks (CNNs) [26, 29, 30], spiking neural networks (SNNs) [31, 32], etc., have been reported, demonstrating remarkable performance on machine learning tasks. The fundamental operations of neural networks, involving data transfers and tensor operations, are achieved through the combination of passive optical devices and high-performance active photonic-electronic components[33, 34]. Specifically, optical signals can be modulated by electrical signals and “multiplied” in accordance with the transmission function of the PIC. The hybrid photonic-electronic platform combines the adaptability of electronic control with the high-speed capabilities of optical computing. Recently, cutting-edge optical processing units have been reported with a matrix processing speed of 3.8 trillion operations per second (TOPS) via time-wavelength multiplexing[30], while other works demonstrated ultra-low power consumption on the order of sub-femtojoules per bit[22].
@@
While integrated photonics offers new opportunities, existing photonic-electronic computing systems still encounter several practical challenges, such as:
@@
  • The inherent challenges in PICs, such as training algorithms, on-chip implementation of nonlinearity for ONNs and system robustness against noise and crosstalk, require careful consideration in both hardware design and software coordination.
This review focuses on recent progress in photonic-electronic integrated circuits for computing. Spanning from digital computing to analog AI accelerators, this paper is structured as follows. Section II begins with an overview of the fundamental blocks in PIC-based digital computing, followed by a survey of recent highlights ranging from the implementation of logic gates to fully functional photonic processing units. Section III focuses on the implementation of ONNs, covering aspects of photonic tensor cores, nonlinearity, and hardware-aware training strategies. Beyond a review from the device and circuit level, Sections IV and V provide a comprehensive analysis of recent photonic AI efforts from the perspectives of accelerator architectures and software-hardware co-design, respectively. The review culminates with Section VI, which offers an outlook on PIC-based optical computing and provides a summarizing conclusion.
II Survey of Optical Digital Computing with PICs
@@ II-A Optical Logic Gates
In the digital domain, both input and output are binary, and the resolution is defined by the number of bits and remains unaffected by the circuit size. A range of building blocks for optical digital computing on integrated photonic platforms, such as optical switches [35], modulators [22, 36], interconnects and photodetectors [37, 38, 39, 40], have been experimentally demonstrated. Basic logic operations (NOT, AND, OR, XOR, etc.), which are fundamental elements of digital systems, have been implemented by diverse PICs. Among these, electro-optic logic, also known as optical-directed logic, has been widely investigated by many research groups as well as foundries. As shown in Fig. 1.a, all E-O devices in the functional block, such as Mach-Zehnder Interferometers (MZIs), microring resonators (MRRs), microdisks, etc., are simultaneously configured by electrical signals. When light traverses the block, optical signals are modulated to execute logic operations in accordance with the PIC design and then propagated downstream or detected by monitors to read out the results. An important feature of the E-O digital logic is that each device is controlled by independent electrical input simultaneously and that signals are transmitted via light without the limitation of RC time constant and delay accumulations inherent in electrical systems. In other words, E-O logic merges the convenience and flexibility of electrical control with the high-speed capabilities of optical computing. Fig. 1.b-c show examples of E-O logic gates using two cascaded MRRs to perform 2-input AND/NAND and OR/NOR operations [41]. The proposed optical logic gate leverages the transmission characteristic of add-drop MRRs, which work as optical switches to implement logic operations and generate complementary outputs at the through and drop ports. Additionally, by configuring the resonant states of MRRs with ‘0’ operands via selecting continuous wave (CW) inputs at either on-resonance or off-resonance wavelengths, a single photonic circuit can execute a variety of logic functions. With consistent logic but varying PIC topologies, XOR and XNOR gates featuring a crossbar structure also have been demonstrated [42] (Fig. 1.d). Operating with similar mechanisms, MZI-based optical gates have also been extensively developed [43, 44].
@@
Figure 1: Implementations of electro-optic logic gates. (a) Schematic diagram of E-O logic composed of passive and active optical components. During each clock cycle, electrical inputs are used to configure the logic circuit, while light carries out logic operations based on the transmission characteristic of the logic block. (b)-(d) Schematic of MRR-based AND/NAND, OR/NOR, and XOR/XNOR gates, as proposed in Ref [41, 42]. In the diagram, the dotted line and the solid line represent MRRs functioning in the “block/pass” and “pass/block” modes, respectively. These configurations correspond to the outputs of ‘0’ and ‘1’ at the through port, given a logic ‘0’ as the electrical input.
@@
In addition to E-O modulation, all-optical logic gate devices have also attracted attention. These devices have been experimentally demonstrated using various structures and phenomena, including photonic crystals [45, 46], surface plasmon polaritons (SPPs) [47], nanowire networks [48], and slot waveguides [49, 50]. A comprehensive classification, explanation of mechanisms, and comparative analysis of all-optical logic gates has been detailed in the prior review [51, 52, 53]. Compared to E-O logic, all-optical logic offers the potential for higher operation speed and bandwidth without extra energy consumption associated with O-E-O conversion. However, the implementation of all-optical logic confronts several practical challenges. Firstly, the complexity and requirements of design and fabrication (e.g., the transmission characteristic of photonic crystal is highly sensitive to its lattice constant) lead to inherent instability and a low contrast ratio between logical states [50]. Secondly, while all-optical gates can be more energy-efficient in signal processing, they often require higher optical signal power (mW-level) to compensate for higher losses or to induce the necessary nonlinear effects for switching. Furthermore, the limited functionality and scalability of all-optical logic, coupled with its higher cost, restrict its widespread application compared to E-O logic.
In recent years, integrated photonics have expanded beyond traditional classical optics, emerging as a compelling platform for quantum information science. Quantum logic gates based on the aforementioned active/passive devices have been widely reported [54, 55, 56]. Quantum PICs offer several significant advantages over bulk optics in the realm of quantum computing. First, PICs enable precise control of phase, polarization, and spatial mode with higher stability, which are essential for manipulating quantum states. Second, silicon exhibits a high third-order nonlinear coefficient $\chi^{(3)}$, facilitates the effective implementation of on-chip single/entangled photon sources through optical processes such as four-wave mixing (FWM) [57]. Third, PICs can integrate the fundamental building blocks of quantum computing—such as photon source, modulators, and single-photon detectors, etc., —in monolithic, hybrid, or heterogeneous configurations [58]. This integration yields scalable, robust, and reconfigurable circuits capable of handling complex quantum computing tasks [59, 57, 60].
II-B Combinational Logic and Reconfigurable PICs
@@
In digital circuits, the output of a combinational logic unit is determined solely by the current input combination, without dependence on previous states. Similarly, the implementation of optical combinational logic could begin with extracting the logical expression from its truth table, followed by designing the corresponding PIC based on the simplified expression. The implementation can rely on assembling fundamental logic gates or leveraging the unique characteristics of optical components or multiplexing techniques for fewer devices and compact layouts. Employing these strategies, diverse optical combinational logic units, including but not limited to adders [61], comparators [62], encoders [63, 64] and decoders [65], have been reported.
@@
The aforementioned optical logic gates and units, tailored for specific tasks, are constrained by a fixed or limited logic representation. The inherent limitation not only complicates the development process but also increases the cost. To address this challenge, reconfigurable PICs offer a promising solution by programming the operational states of optical switches within a pre-designed framework using additional signals. The input operands and the reconfiguration signal can be independently managed by two separate control units within the modulator, such as the dual arms of MZIs. Generally, reconfiguration signals do not require high-speed modulation to the same extent as the input signals used as logic operands. Qiu et al. proposed a reconfigurable logic unit based on MRRs embedded with two modulation mechanisms [66]. As is shown in Fig. 2.a, the logic operand is modulated at high speeds via the p-i-n junction operating in the carrier injection mode, while the resonant state of MRR can be reconfigured using the microheater. Additionally, multi-operand modulators and non-volatile devices are also promising candidates for reconfigurable PICs [67, 68, 69].
@@
$\vspace{-5pt}Y=y_{1}+y_{2}+\ldots+y_{m},\text{ where }y_{i}=\prod_{k=1}^{n}X_{k}\text{ or }\overline{X_{k}}\vspace{-3pt}$ (1)
@@
In this expression, $\overline{X_{k}}$ denotes the complement logic of input $X_{k}$. Using reconfigurable optical switches, the architecture illustrated in Fig. 2.b-c theoretically can implement arbitrary logic functions conforming to the expression format presented in Eq. (1). The product term $y_{i}$, i.e., the logic AND operation, can be implemented by $n$ serially connected reconfigurable optical switches along single bus waveguides. This block yields logic ‘1’ output only when all switches are in the “pass” state. Reconfiguration signals $R_{i}$ are used to determine whether the corresponding operand contributes complement logic to $y_{i}$. In contrast to electrical digital computing, the OR operation for optical signals can be directly implemented using a combiner and detected by a photodetector. It is important to recognize that when product terms are represented by the same wavelength, the amalgamation may result in logical errors due to coherent interference. This issue can be avoided either by using distinct wavelengths for each branch [66], or equipping each branch with photodetectors individually [70]. Besides, given that photodetectors operate as current sources, a straightforward parallel configuration could achieve electrical OR logic without introducing additional delays.
@@
Figure 2: Reconfigurable PICs for arbitrary combinational logic. (a). Optical micrograph and cross-sectional diagram of a reconfigurable MRR featuring two modulation mechanisms. The p-i-n junction is applied by RF signal for input encoding, while the microheater is connected to the low-speed DC signal for resonance mode reconfiguration [66]. (b)-(c) Schematic of a PIC architecture that enables the implementation of arbitrary combinational logic expressions based on reconfigurable optical switchs (ROSs).
While this strategy offers a general solution for optical combinational logic, the limited scalability restricts its practical applicability in scenarios involving numerous operands. For an $n$-operands system, the complexity of this processing unit escalates as $O(n\cdot 2^{n})$, indicating an exponential increase in the requisite number of switches with $n$. Furthermore, issues related to power consumption and accumulated losses also need to be considered. The dynamic power of the system $P_{\text{dynamic}}$ can be expressed as:
$P_{\text{dynamic}}=\frac{1}{4}\alpha CV^{2}f$ (2)
where, $\alpha$ is the activity factor and $C$ is the total capacitance of the E-O modulators, which increases proportionally with the number of modulators. It is important to note that while Eq. (2) has a similar expression to that used for CMOS transistors, the supply voltage $V$ does not necessarily scale with $f$. Therefore, to reduce $P_{\text{dynamic}}$, two straightforward strategies can be considered: (1) Decrease the number of modulators, for example, by employing multi-operand logic gates to squeeze logic functions into fewer devices [68]; (2) Utilize devices with low capacitance, such as microdisks with capacitance in 10’s fF [24]. Additionally, when present, thermal tuning typically dominates the power consumption for modulation. This portion of power consumption can be reduced or even eliminated through device-level optimization, such as post-fabrication trimming [71], or using energy-efficient tuning mechanisms (discussed in Section III-A). The total power consumption also includes the laser and detection parts, as well as the potential static power consumption. The laser power is primarily determined by the losses, typically at the mW level with dB/cm-level loss, which also correlates with the PIC scale.
@@
In the realm of electrical digital design, electronic design automation (EDA) tools assist designers by automatically generating and optimizing circuits from logic expressions, as well as auto-placing and routing in physical design. With the advancement of PIC-based computing, logic synthesis methods specified for PIC design have been proposed, enabling large-scale design automation and optimization of compact PICs for digital computing [72, 73, 74].
II-C Toward a Fully-functional EPALU